In a regular polygon, the measure of each angle is 144 degree. How many sides does the polygon have?: In a regular polygon, the measure of each angle is 144 degree. How many sides does the polygon have?
Answer by Earlsdon(3717)

In a regular polygon of n sides, the measure of each of the (equal) interior angles is given by:
A[i] = (n-2)180/n Substitute A[i] = 144 and solve for n.
144 = (n-2)180/n Multiply both sides by n and simplify.
144n = 180n-360 Add 360 to both sides and subtract 144n from both sides.
360 = 36n Divide both sides by 36.
n = 10 A decagon.